OpenStudy (anonymous):

Find all values of x (if any) where the tangent line to the graph of the given equation is horizontal f(x)=-4^2-3x not so much looking for a direct answer just wondering how i would solve this question

OpenStudy (anonymous):

You'd start by finding the derivative of the equation first

OpenStudy (anonymous):

Yes, the solutions to the first derivative are the points where are max/min (horizontal tangent lines) of the parent function.

OpenStudy (anonymous):

Assuming this is a calculus question

OpenStudy (anonymous):

okay the derivative is -3 then?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thats my answer?

OpenStudy (anonymous):

Is there supposed to be an x in -4^2?

OpenStudy (anonymous):

oopes i meant -8*x-3 is the derivative

OpenStudy (anonymous):

What is the given equation, exactly?

OpenStudy (anonymous):

f(x)=-4x^2-3x my bad

OpenStudy (anonymous):

but the derivative of that is -8x-3

OpenStudy (anonymous):

Then yes the answer is the solutions to -8x-3

OpenStudy (anonymous):

but how would i find any solutions if i wasnt given any more information?

OpenStudy (anonymous):

How much less can they provide? They've given you the equation, unless they want you to perform a proof with just variables maybe

OpenStudy (anonymous):

i even entered none and it said incorrect sooo im really not sure

OpenStudy (anonymous):

You need to find the vertex of this parabola, because that is the point where the slope will be 0.

OpenStudy (anonymous):

how would i find the vertex?

OpenStudy (anonymous):

I'd put the parabola in the standard form to identify the vertex easier. This one seems relatively easy to complete the square at first glance, so I'd give it a shot. Do you know how to do that?

OpenStudy (anonymous):

no i dont:(

OpenStudy (anonymous):

You have to factor the coefficient of the leading term first \[-4(x^2 + \frac{3}{4}x)\] \[-4[(x^2 + \frac{3}{4}x + \frac{9}{64}) - \frac{9}{64}] = -4[(x + \frac{3}{8})^2 - \frac{9}{64}] \] distribute to expand and get the standard form \[-4(x + \frac{3}{8})^2 + \frac{9}{16}\]So the vertex is (-3/8, 9/16)

OpenStudy (anonymous):

so after we have found the vertex what do we do next?